Mary,aldo,and Lamar sent a total of 99 text messages over their cell phones during the weekend. Aldo sent 9 fewer messages than Mary. Lamar sent 4 times as many messages as Aldo. How many messages did they each send?
Answer by ikleyn(52803)

.

Mary x

Aldo (x-9)

Lamar 4*(x-9)


    x + (x-9) + 4*(x-9) = 99

    6x - 45             = 99

    6x                  = 99 + 45

    6x                  = 144

     x                  = 144/6 = 24.


From this point, complete the solution on your own.

You obtained all needed instructions.